Consider an isolated system, consisting of a box of volume V0 with a
partition which separates it into two subsystems; the first subsystem has
a volume of 0.4 V0 and contains one mole of an ideal gas at 300 K, while
the other is evacuated. Calculate the entropy change in the gas during
the process of its (free) expansion on removing the partition.
1
Expert's answer
2020-12-28T10:28:31-0500

Answer

Change in entropy can be expressed as

"\Delta S=nR\ln(\frac{v_f}{v_i})"

"\Delta S=1\times8.314\times\ln(\frac{1}{0.4})"

"\Delta S=1\times8.314\times0.92"

"\Delta S=7.65J/K"